OpenStudy (aaronandyson):

Prove That (cosecA-sinA)(secA-cosA)(tanA+cotA) = 2

OpenStudy (aaronandyson):

@k_lynn

OpenStudy (k_lynn):

Sorry, I"m not very good at that kind of math.

OpenStudy (mathstudent55):

\((\csc A- \sin A)(\sec A- \cos A)(\tan A+ \cot A) = 2\) \((\dfrac{1}{\sin A} - \sin A)(\dfrac{1}{\cos A}- \cos A)(\dfrac{\sin A}{\cos A}+ \dfrac{\cos A}{\sin A}) \) \((\dfrac{1}{\sin A} - \sin A \times \dfrac{\sin A}{\sin A})(\dfrac{1}{\cos A}- \cos A \times \dfrac{\cos A}{\cos A})(\dfrac{\sin A}{\cos A} \times \dfrac{\sin A}{\sin A}+ \dfrac{\cos A}{\sin A}\times \dfrac{\cos A}{\cos A}) \) Can you continue?

OpenStudy (mathstudent55):

Sorry, the last line above was cut off. Here it is: \((\dfrac{1}{\sin A} - \sin A \times \dfrac{\sin A}{\sin A})(\dfrac{1}{\cos A}- \cos A \times \dfrac{\cos A}{\cos A}) \) \( \times (\dfrac{\sin A}{\cos A} \times \dfrac{\sin A}{\sin A}+ \dfrac{\cos A}{\sin A}\times \dfrac{\cos A}{\cos A}) \)

OpenStudy (mathstudent55):

In each set of parentheses, add or subtract the fractions using the common denominators written. Then use trig identity substitutions.

OpenStudy (aaronandyson):

??????????

OpenStudy (aaronandyson):

What???

OpenStudy (aaronandyson):

@mathmate

OpenStudy (aaronandyson):

@rvc

OpenStudy (aaronandyson):

@satellite73

OpenStudy (aaronandyson):

@jigglypuff314

OpenStudy (mathstudent55):

\((\dfrac{1}{\sin A} - \sin A \times \dfrac{\sin A}{\sin A})(\dfrac{1}{\cos A}- \cos A \times \dfrac{\cos A}{\cos A})\) \(\times (\dfrac{\sin A}{\cos A} \times \dfrac{\sin A}{\sin A}+ \dfrac{\cos A}{\sin A}\times \dfrac{\cos A}{\cos A})\) \((\dfrac{1}{\sin A} - \dfrac{\sin^2 A}{\sin A})(\dfrac{1}{\cos A}- \dfrac{\cos^2 A}{\cos A}) (\dfrac{\sin^2 A}{\cos A \sin A} + \dfrac{\cos^2 A}{\cos A\sin A})\) \((\dfrac{1 -\sin^2 A}{\sin A})(\dfrac{1 - \cos^2 A}{\cos A}) (\dfrac{\sin^2 A + \cos^2 A}{\cos A\sin A})\) \((\dfrac{\cos^2 A}{\sin A})(\dfrac{\sin^2 A}{\cos A}) (\dfrac{1}{\cos A\sin A})\) \(1\) I get 1 on the left side, not 2.